OpenStudy (anonymous):

collect like terms 14x+7y-11x-9y=

OpenStudy (angela210793):

3x-2y

OpenStudy (anonymous):

anglea i came up with an answer 3 only

OpenStudy (angela210793):

3? well you add them as real no.s (is just tht they have a x and y behind) 14-11=3 right? so 3x 7-9=-2 so 2y 14x+7y-11x-9y=3x-2y

OpenStudy (anonymous):

gotcha thanks

OpenStudy (angela210793):

uw :)

OpenStudy (anonymous):

can you help with this one graph the system of inequalities y>=to-5 x>=to6

OpenStudy (angela210793):

well...wolframalpha will....w8

OpenStudy (anonymous):

so angela i would shade this graph in correct

OpenStudy (angela210793):

yea

OpenStudy (anonymous):

k